# Frequency Diverse Array OFDM System for Joint Communication and Sensing

### v1

@@ II-A Frequency Diverse Array
where $\lambda_{\mathrm{c}}=c/f_{c}$. The array factor of the FDA without tapering is then given by
$\mathrm{AF_{FDA}}(\theta,t)$ $=\sum_{m=-(M-1)/2}^{(M-1)/2}{e^{j2\pi m\left(\Delta ft+\frac{d_{\mathrm{a}}}{\lambda_{\mathrm{c}}}\sin{(\theta)}\right)}}$
$=\frac{\sin{\left(M\pi\left(\Delta ft+\frac{d_{\mathrm{a}}}{\lambda_{\mathrm{c}}}\sin{(\theta)}\right)\right)}}{\sin{\left(\pi\left(\Delta ft+\frac{d_{\mathrm{a}}}{\lambda_{\mathrm{c}}}\sin{(\theta)}\right)\right)}}.$ (5)
@@
Figure 2: Evolution of the FDA beam pattern and gain [dB] over time for ULA with $M=8$ antennas and $\Delta f=1$ kHz.
@@ IV-B Frequency shift
The signal transmitted by the $m$th antenna is a frequency-shifted version of the original signal
$y_{m}(n)=y(n)e^{j2\pi m\frac{\varepsilon}{N}n},$ (15)
where $\varepsilon=\frac{\Delta f}{\Delta f_{\mathrm{sc}}}$ is the frequency offset normalized by the subcarrier spacing. The frequency domain representation of the frequency-shifted signal is obtained by performing a DFT of the same length as in (14). Following the formulation in [12], the signal at antenna $m$ can be expressed as
@@
$=\frac{1}{N}\sum_{l=0}^{N-1}s(l)\underbrace{\sum_{n=0}^{N-1}e^{j2\pi n\frac{(l+m\varepsilon-k)}{N}}}_{\mathrm{ICI}},$ (16)
where $l$ is another subcarrier index. Due to frequency shift the subcarriers are no longer orthogonal and intercarrier interference (ICI) appears. The ICI component between the $l$th and $k$th subcarrier can be written as [12]
@@
$y_{m}(k)$ $=\sum_{l=0}^{N-1}s(l)p_{m}(l-k)$
$=s(k)p_{m}(0)+\sum_{\begin{subarray}{c}l=0\\ l\neq k\end{subarray}}^{N-1}s(l)p_{m}(l-k).$ (18)
@@ IV-D Sensing
The reflection signal from the $m$th antenna is then
@@
$=$ $\gamma e^{-j2\pi lT_{\mathrm{p}}\frac{v}{\lambda_{\mathrm{c}}}}$
$e^{-j2\pi(m\Delta f+k\Delta f_{\mathrm{sc}})\left(\frac{2d}{c}+\frac{md_{\mathrm{a}}}{c}\left(\sin{(\theta_{\mathrm{TX}}(t))}-\sin{(\theta_{q})}\right)\right)}$ (26)
@@
$x_{m}(k,l,t)=\gamma e^{j(\phi_{l}+2\varphi_{k})}e^{-j2\pi k\Delta f_{\mathrm{sc}}\frac{md_{\mathrm{a}}}{c}\Delta\phi_{q}(t)},$ (27)
where $\phi_{l}=-j2\pi lT_{\mathrm{p}}v/\lambda_{\mathrm{c}}$ is the phase shift due to the target velocity and $\Delta\phi_{q}(t)=\sin{(\theta_{\mathrm{TX}}(t))}-\sin{(\theta_{q})}$. The sum of the reflected signals is then

### v2

@@ II-A Frequency Diverse Array
where $\lambda_{\mathrm{c}}=c/f_{c}$. Given that the power is uniformly split across antennas the array factor of the FDA is given by
$\mathrm{AF_{FDA}}(\theta,t)$ $=\frac{1}{\sqrt{M}}\sum_{m=-(M-1)/2}^{(M-1)/2}{e^{j2\pi m\left(\Delta ft+\frac{d_{\mathrm{a}}}{\lambda_{\mathrm{c}}}\sin{(\theta)}\right)}}$
$=\frac{1}{\sqrt{M}}\frac{\sin{\left(M\pi\left(\Delta ft+\frac{d_{\mathrm{a}}}{\lambda_{\mathrm{c}}}\sin{(\theta)}\right)\right)}}{\sin{\left(\pi\left(\Delta ft+\frac{d_{\mathrm{a}}}{\lambda_{\mathrm{c}}}\sin{(\theta)}\right)\right)}}.$ (5)
@@
Figure 2: Evolution of the FDA beam pattern and normalized gain [dB] over time for ULA with $M=8$ antennas and $\Delta f=1$ kHz.
@@ IV-B Frequency shift
The signal transmitted by the $m$th antenna is a scaled and frequency-shifted version of the original signal
$y_{m}(n)=\frac{1}{\sqrt{M}}y(n)e^{j2\pi m\frac{\varepsilon}{N}n},$ (15)
where $\varepsilon=\frac{\Delta f}{\Delta f_{\mathrm{sc}}}$ is the frequency offset normalized by the subcarrier spacing. The factor $1/\sqrt{M}$ normalizes the total transmit power to unity. The frequency domain representation of the frequency-shifted signal is obtained by performing a DFT of the same length as in (14). Following the formulation in [12], the signal at antenna $m$ can be expressed as
@@
$=\frac{1}{N\sqrt{M}}\sum_{l=0}^{N-1}s(l)\underbrace{\sum_{n=0}^{N-1}e^{j2\pi n\frac{(l+m\varepsilon-k)}{N}}}_{\mathrm{ICI}},$ (16)
where $l$ is the other subcarrier index. Due to frequency shift the subcarriers are no longer orthogonal and intercarrier interference (ICI) appears. The ICI component between the $l$th and $k$th subcarrier can be written as [12]
@@
$y_{m}(k)$ $=\frac{1}{\sqrt{M}}\sum_{l=0}^{N-1}s(l)p_{m}(l-k)$
$=\frac{1}{\sqrt{M}}\left(s(k)p_{m}(0)+\sum_{\begin{subarray}{c}l=0\\ l\neq k\end{subarray}}^{N-1}s(l)p_{m}(l-k)\right).$ (18)
@@ IV-D Sensing
The reflection signal from the $m$th antenna is
@@
$=$ $\frac{1}{\sqrt{M}}\gamma e^{-j2\pi lT_{\mathrm{p}}\frac{v}{\lambda_{\mathrm{c}}}}$ (26)
$e^{-j2\pi(m\Delta f+k\Delta f_{\mathrm{sc}})\left(\frac{2d}{c}+\frac{md_{\mathrm{a}}}{c}\left(\sin{(\theta_{\mathrm{TX}}(t))}-\sin{(\theta_{q})}\right)\right)}$
@@
$x_{m}(k,l,t)=\frac{1}{\sqrt{M}}\gamma e^{j(\phi_{l}+2\varphi_{k})}e^{-j2\pi k\Delta f_{\mathrm{sc}}\frac{md_{\mathrm{a}}}{c}\Delta\phi_{q}(t)},$ (27)
where $\phi_{l}=-j2\pi lT_{\mathrm{p}}v/\lambda_{\mathrm{c}}$ is the phase shift due to the target velocity and $\Delta\phi_{q}(t)=\sin{(\theta_{\mathrm{TX}}(t))}-\sin{(\theta_{q})}$. The sum of the reflected signals is